solve 4(x-2)^2=36 by using square roots.
Answer by lwsshak3(11628)

solve4(x -2)^2=36 by using square roots.
4(x -2)^2=36
divide by 4
(x-2)^2=9
take the sqrt of both sides
x-2=±√9=±3 (note: in an equation, the sqrt of a number always has two roots, one plus and one minus. When not in an equation, by itself, the sqrt has one positive root.)
x=2±3
x=-1
or
x=5